# Higgs Phases and Boundary Criticality in memoriam Peter Higgs

### v1

@@ II.1.3 Magnetic Monopoles
II.2 Open Boundaries and Global Symmetry
We address now a well-known, but subtle point: by taking $\lambda$ in (4) to be a constant function, it appears at first sight that this theory has a global 0-form U(1) symmetry, shifting $\theta_{i}\to\theta_{i}+\lambda$, leaving $A$ unchanged since $\mathrm{d}\lambda=0$. In the Hamiltonian picture, this transformation is generated by the total charge,
@@ II.2 Open Boundaries and Global Symmetry
thus such a symmetry corresponds to global conservation of electric charge. Note, however, that in the absence of boundaries this “global symmetry” is pure gauge, because all physical quantum states belong to the zero-charge sector and thus carry the same quantum number. By the Gauss law, Eq. 13, $\hat{Q}_{\text{bulk}}$ is exactly zero for a system with periodic boundary conditions. In other words, since all electric flux lines must end somewhere inside the system, the system must be globally charge neutral. By construction, such a “symmetry” is therefore trivial and cannot be explicitly or spontaneously broken. However, in presence of specific boundary conditions, these global U(1) transformations actually generate a physical global symmetry that acts on the boundaries which can be spontaneously broken [19, 20].
We consider a lattice with open boundaries in the form illustrated in Fig. 2. The bulk of the system is a (hyper)cubic lattice with sites indicated by white spheres, links by black lines, and plaquettes by gray faces. At the boundary, we include a layer of cubic cells (green) which separate the bulk from the vacuum (blue sites, edges, and plaquettes). In particular, there is a set of links bridging between the bulk and the vacuum which carry dynamical gauge degrees of freedom and can therefore support electric flux lines which effectively exit the system. The vacuum side (blue) does not contain any dynamical degrees of freedom.
Key to our choice of boundary conditions is that we demand that the Gauss law, Eq. 13, is respected at every bulk site (white sphere), including those at the ends of the boundary links. No Gauss law constraints are imposed at vacuum sites. Let $A_{i}$ denote the gauge potential on the boundary link touching site $i$, oriented “in” from the vacuum to the bulk. For the Gauss law to be respected at $i$, we must have that under the gauge transformation Eq. 4, $A_{i}\to A_{i}+\lambda_{i}$, i.e. it is “uncompensated” at the vacuum end of the link.
@@
$\hat{W}[\gamma_{\text{open}}]=\exp\left(i\left[\hat{\theta}_{i}-\sum_{\mathclap{\ell\in\gamma_{\text{open}}}}\hat{A}_{\ell}\right]\right),$ (16)
where $\gamma_{\text{open}}$ is a contour starting in the vacuum and ending at bulk site $i$. These operators create an isolated charge in the bulk, attached to an electric flux line that exits through the boundary, which is not possible in a closed system. These half-open string operators are charged under the global transformations generated by $\hat{Q}_{\text{bulk}}$, Eq. 15. By the Gauss law, $\hat{Q}_{\text{bulk}}$ is equivalent to the net electric flux through the boundary,
@@
where the sum is over all boundary links (green in Fig. 2) oriented out. If the Hamiltonian contains no half-open string operators, then the bulk charge is conserved, and there is a global U(1) symmetry. Eq. 17 defines a “bulk-boundary correspondence” between charge and flux, and generates a global symmetry which may either be seen as acting on the bulk matter degrees of freedom or on the boundary gauge degrees of freedom.
@@
Figure 3: Boundary criticality for the 4D $U(1)$ Higgs phase: (a) Average boundary plaquette for $\kappa=$0.55, 0.64, 0.85, 0.94, 1.0 for $L=16$. The transition shifts towards smaller $\beta$ for larger values of $\kappa$. (b) The Binder ratio $U_{4}$ for the magnetization at the boundary as a function of $\beta$ for $\kappa=2$ and $L=16,20,24,28,32$. (c) Rescaled Binder ratio showing collapse for $\nu=0.67$ corresponding to the 3D XY universality class.
II.3 Boundary Symmetry Breaking in the Abelian Higgs Model
The theory can now be chosen such that the Hamiltonian commutes with the charge $\hat{Q}_{\text{bulk}}$. This now-physical global $U(1)$ symmetry corresponds to global charge conservation or, equivalently, conservation of flux through the boundary. Since the open system has a global symmetry on the boundary, it may be spontaneously broken, a scenario we now study. With the boundary conditions shown in Fig. 2, the Euclidean action that we will study is given by
@@ II.3 Boundary Symmetry Breaking in the Abelian Higgs Model
with the boundary portion given by the Wilson plaquette loops on the boundary plaquettes (light green in Fig. 2),
@@
II.3.1 Boundary XY Model at Infinite $\kappa$
To begin, we consider the behavior of this theory in the $\kappa\to\infty$ limit, i.e. deep in the Higgs regime. From the bulk action, Eq. 8, in this limit the bulk satisfies the constraint $A=\mathrm{d}\theta$, i.e. the bulk gauge field is exact and thus pure gauge. Indeed there are no physical degrees of freedom left in the bulk—rotating to unitary gauge, $\theta=\text{const}.$, we end up with $A=0$ on all bulk links and a vanishing matter field. However, no such constraint is enforced on the boundary links bridging between the bulk and vacuum, and the gauge field on these links is free to fluctuate. Thus we obtain a dynamical 3D theory on the boundary of the system governed by the boundary action in Eq. 19.
In this limit, referring to Fig. 2, each boundary plaquette (green) has one edge in the bulk (black) with $A_{ij}=\theta_{j}-\theta_{i}$, and two edges straddling between the bulk and vacuum (green) which remain dynamical degrees of freedom. Substituting this into Eq. 20, we can recombine terms into the gauge-invariant variables
@@ II.3.1 Boundary XY Model at Infinite $\kappa$
$S^{\text{bdry}}_{\kappa\to\infty}(\beta)=-\beta\sum_{\mathclap{\langle ij\rangle\in\partial X}}\cos(\vartheta_{j}-\vartheta_{i}),$ (22)
which is a 3D XY model at inverse temperature $\beta$. This must exhibit a continuous phase transition from a paramagnet at small $\beta$ to a spontaneously broken phase at large $\beta$. Thus we infer that along the $\kappa=\infty$ line in the phase diagram there is a boundary phase transition in the 3D XY universality class indicated at the top of Fig. 1. We note that in the case of the $\mathbbm{Z}_{2}$ Abelian-Higgs model, the same mechanism generates a boundary Ising model at $\kappa=\infty$ [19].
II.3.2 Boundary Phase Transition at Finite $\kappa$
@@
These arguments are suggestive of the picture illustrated in Fig. 1, with a boundary phase transition between the Higgs and confinement regimes of the bulk phase diagram. To test this assertion, we have carried out Monte Carlo simulations of the full 4D lattice gauge theory with boundary. We compute the local XY order parameter $\langle\vartheta_{i}\rangle$ on the boundary as well as gauge invariant bulk observables $\langle\Lambda_{\ell}\rangle$ and $\langle W_{p}\rangle$. The results are summarized in Fig. 3. We find clear signs of a boundary phase transition, with Fig. 3(a) showing the boundary order parameter behavior as a function of $\beta$ while holding $\kappa$ fixed showing behavior consistent with a continuous boundary phase transition. Fig. 3(b) shows the Binder cumulant for the order parameter taken along a cut at $\kappa=2$ for different system sizes ranging from $L=16$ to $L=32$, showing crossing behavior consistent with a second-order transition. Lastly, Fig. 3(c) shows the Binder parameter with $\beta$ scaled by $L^{1/\nu}$ using the 3D XY critical exponent $\nu\approx 0.67$ [31], showing excellent scaling collapse, confirming a second-order phase transition on the boundary even for only moderately large $\kappa$.
Monte Carlo simulations of the 3D XY model in the literature put the critical point at $\beta_{c}=0.45420(2)$ [31]. We find that $\beta_{c}$ tends towards this value in the large $\kappa$ limit. On general grounds we should expect that bulk fluctuations will serve to disorder the boundary. Therefore, by lowering $\kappa$ we expect the transition shifts to larger $\beta$ (lower effective temperature in the statistical model). This is indeed what we observe numerically. For even smaller $\kappa$ we find that the boundary transition line appears to intercept the bulk critical endpoint, as illustrated in Fig. 1. Below, by changing the boundary theory, we demonstrate that this result is not universal.
II.3.3 Tuning the Boundary Coupling
We now consider tuning the boundary coupling in Eq. 19 relative to the bulk, parameterized by the dimensionless ratio
@@ II.3.3 Tuning the Boundary Coupling
Indeed, this is precisely what we find numerically with the resulting transition lines shown in Fig. 4. We find, for all $\alpha$, that the transition is present and that it drifts to larger $\beta$ as $\kappa$ is reduced, consistent with having to lower the temperature more to suppress the enhanced matter field fluctuations. For $\alpha<1$ the line shifts to larger values of $\beta$, and appears to separate from the tricritical point, instead terminating on the first order transition line separating the Higgs phase from the deconfined phase. For $\alpha>1$, the boundary transition line shifts to smaller values of $\beta$, and appears to continue to terminate on the tricritical point.
@@
II.4 Summary and Discussion: Abelian Case
In this section we have explored the case of the U(1) Abelian-Higgs model in $D$ spacetime dimensions with open boundary conditions of the form shown in Fig. 2, governed by the action defined by Eqs. 8, 19 and 18. In the case of open boundaries there is a global bulk charge symmetry for this model, because charge is not allowed to enter or leave the system. By the Gauss law, this is physically equivalent to a symmetry acting on the boundary of the system, yielding conservation of total electric flux through the boundary. We investigated the possibility of spontaneously breaking this symmetry.
In the $\cramped{\kappa\to\infty}$ limit, the bulk degrees of freedom are fully frozen while the boundary degrees of freedom remain fluctuating. The boundary theory in this limit can be written in the gauge-invariant form of a $(D-1)$-dimensional 0-form U(1) model, i.e. an XY model, Eq. 22, which therefore exhibits a boundary phase transition at a critical value of $\beta$ in the appropriate XY universality class. By performing explicit numerical simulation in $D=4$ Euclidean spacetime dimensions and measuring the gauge-invariant order parameter, Fig. 3, we found that this boundary phase transition persists away from the $\cramped{\kappa\to\infty}$ limit, and appears to stay in the 3D XY universality class. The transition line appears to terminate at a critical point in the bulk phase diagram and can be tuned by adjusting the boundary coupling relative to the bulk coupling, as shown in Fig. 4.
This boundary transition line appears to delineate between the Higgs and confining regimes of the phase diagram. The idea that the Higgs phase can be characterized by boundary symmetry breaking was first raised in [19, 20]. This gives physical meaning to the non-gauge-invariant adage that the Higgs regime is a charge condensate by imposing open boundaries that make the symmetry physical. We note, however, that by tuning the boundary coupling relative to the bulk coupling, the location of the transition line in the phase diagram can be moved. This raises questions as to what the boundary phase transition implies as a probe of bulk physics, which we defer to the discussion at the end of the paper (Section V.2). First we broaden our perspective by studying more complex non-Abelian models exhibiting similar physics in Section III.
A few comments on extensions and exceptional cases are in order before we proceed. These results naturally carry over to the discrete Abelian gauge groups $\mathbbm{Z}_{N}$ by restricting the Higgs field $\theta_{i}$ and gauge field $A_{ij}$ to take discrete values in multiples of $2\pi/N$. The boundary theory Eq. 22 then becomes an $N$-state clock model. The $\mathbbm{Z}_{2}$ case was discussed in [19] where the boundary theory was shown to be an Ising model. They also can be generalized to higher-form extensions of the Abelian-Higgs model, which we discuss further in Section IV.
@@ II.4 Summary and Discussion: Abelian Case
III Boundary Symmetry Breaking in Non-Abelian Higgs Models
@@
III.1 Preliminaries: Lattice Formulation
@@
where $g_{i}\in\mathcal{G}$ are arbitrary group elements associated to each site $i$. The minimal gauge-invariant quantity is the Wilson plaquette-loop (compare to Eq. 6),
$W_{p}=\frac{1}{\mathrm{dim}({r})}\,\mathrm{Tr}_{\,{r}}\,\,{\prod_{\mathclap{\ell\in\partial p}}}^{P}U_{\ell}$ (27)
where the superscript $P$ on the product indicates path-ordering, and the trace may be taken in a representation ${r}$. Normalizing by the dimension of the representation ensures that the trivial Wilson loop has unit magnitude. We focus primarily on the cases $\mathrm{SU}(N)$ and $\mathrm{SO}(N)$, taking the trace in the fundamental representation as $N\times N$ matrices.
III.1.1 Higgs Fields
For the Higgs field, many different models can be considered by putting the Higgs field in different representations of the gauge group. We consider two different types here for concreteness: vector-valued (fundamental representation) Higgs, and group-valued Higgs. These are both possible extensions of the Abelian U(1) rotor model considered in Section II, since a rotor may be viewed either as a fixed-length vector, or as a U(1) group element. In either case, the action is given by the Fradkin-Shenker form, Eq. 7, the only difference being the definition of the Wilson link $\Lambda_{\ell}$. The generating function for the quantum theory is given by the Euclidean path integral, where the integration over the group-valued variables is performed with respect to the Haar measure.
The familiar model is the fundamental-Higgs, where the Higgs field is an $N$-component vector, as in the Standard Model and analogous to Eq. 1. Denoting the Higgs vector at site $i$ by $\phi_{i}\in V_{i}$, we freeze the radial mode as in the Abelian Fradkin-Shenker model. Gauge transformations rotate the Higgs field as $\phi_{i}\to g_{i}^{\text{f}}\phi_{i}$ where $g_{i}^{\text{f}}$ is a group element in the fundamental matrix representation. The group-valued link variables define parallel-transport maps for the Higgs field, $U_{ij}:V_{j}\to V_{i}$, i.e. they related the color frames at neighboring sites, and the generalization of the gauge-invariant Wilson link observable, Eq. 5, is

### v2

@@ 2.1.3 Magnetic Monopoles
2.2 Open Boundaries and Global Symmetry
We address now a well-known, but subtle point: by taking $\lambda$ in Eq.˜4 to be a constant function, it appears at first sight that this theory has a global 0-form U(1) symmetry, shifting $\theta_{i}\to\theta_{i}+\lambda$, leaving $A$ unchanged since $\mathrm{d}\lambda=0$. In the Hamiltonian picture, this transformation is generated by the total charge,
@@ 2.2 Open Boundaries and Global Symmetry
thus such a symmetry corresponds to global conservation of electric charge. Note, however, that in the absence of boundaries this “global symmetry” is pure gauge, because all physical quantum states belong to the zero-charge sector and thus carry the same quantum number. By the Gauss law, Eq.˜13, $\hat{Q}_{\text{bulk}}$ is exactly zero for a system with periodic boundary conditions. In other words, since all electric flux lines must end somewhere inside the system, the system must be globally charge neutral. By construction, such a “symmetry” is therefore trivial and cannot be explicitly or spontaneously broken. However, in presence of specific boundary conditions, these global U(1) transformations actually generate a physical global symmetry that acts on the boundaries which can be spontaneously broken [22, 23].
We consider a lattice with open boundaries in the form illustrated in Fig.˜2. The bulk of the system is a (hyper)cubic lattice with sites indicated by white spheres, links by black lines, and plaquettes by gray faces. At the boundary, we include a layer of cubic cells (green) which separate the bulk from the vacuum (blue sites, edges, and plaquettes). In particular, there is a set of links bridging between the bulk and the vacuum which carry dynamical gauge degrees of freedom and can therefore support electric flux lines which effectively exit the system. The vacuum side (blue) does not contain any dynamical degrees of freedom.
Key to our choice of boundary conditions is that we demand that the Gauss law, Eq.˜13, is respected at every bulk site (white sphere), including those at the ends of the boundary links. No Gauss law constraints are imposed at vacuum sites. Let $A_{i}$ denote the gauge potential on the boundary link touching site $i$, oriented “in” from the vacuum to the bulk. For the Gauss law to be respected at $i$, we must have that under the gauge transformation Eq.˜4, $A_{i}\to A_{i}+\lambda_{i}$, i.e. it is “uncompensated” at the vacuum end of the link.
@@
[…]
where $\gamma_{\text{open}}$ is a contour starting in the vacuum and ending at bulk site $i$. These operators create an isolated charge in the bulk, attached to an electric flux line that exits through the boundary, which is not possible in a closed system. These half-open string operators are charged under the global transformations generated by $\hat{Q}_{\text{bulk}}$, Eq.˜15. By the Gauss law, $\hat{Q}_{\text{bulk}}$ is equivalent to the net electric flux through the boundary,
@@
where the sum is over all boundary links (green in Fig.˜2) oriented out. If the Hamiltonian contains no half-open string operators, then the bulk charge is conserved, and there is a global U(1) symmetry. Eq.˜17 defines a “bulk-boundary correspondence” between charge and flux, and generates a global symmetry which may either be seen as acting on the bulk matter degrees of freedom or on the boundary gauge degrees of freedom.
@@
Figure 3: Boundary criticality for the 4D $U(1)$ Higgs phase: (a) Average boundary plaquette for $\kappa=$0.55, 0.64, 0.85, 0.94, 1.0 for $L=16$. The transition shifts towards smaller $\beta$ for larger values of $\kappa$. (b) The Binder ratio $U_{4}$ for the magnetization at the boundary as a function of $\beta$ for $\kappa=2.0$ and $L=16,20,24,28,32$. (c) Rescaled Binder ratio for $\kappa=2.0$ showing collapse for $\nu=0.67$ corresponding to the 3D XY universality class.
2.3 Boundary Symmetry Breaking in the Abelian Higgs Model
The theory can now be chosen such that the Hamiltonian commutes with the charge $\hat{Q}_{\text{bulk}}$. This now-physical global $U(1)$ symmetry corresponds to global charge conservation or, equivalently, conservation of flux through the boundary. Since the open system has a global symmetry on the boundary, it may be spontaneously broken, a scenario we now study. With the boundary conditions shown in Fig.˜2, the Euclidean action that we will study is given by
@@ 2.3 Boundary Symmetry Breaking in the Abelian Higgs Model
with the boundary portion given by the Wilson plaquette loops on the boundary plaquettes (light green in Fig.˜2),
@@
2.3.1 Boundary XY Model at Infinite $\kappa$
To begin, we consider the behavior of this theory in the $\kappa\to\infty$ limit, i.e. deep in the Higgs regime. From the bulk action, Eq.˜8, in this limit the bulk satisfies the constraint $A=\mathrm{d}\theta$, i.e. the bulk gauge field is exact and thus pure gauge. Indeed there are no physical degrees of freedom left in the bulk—rotating to unitary gauge, $\theta=\text{const}.$, we end up with $A=0$ on all bulk links and a vanishing matter field. However, no such constraint is enforced on the boundary links bridging between the bulk and vacuum, and the gauge field on these links is free to fluctuate. Thus we obtain a dynamical 3D theory on the boundary of the system governed by the boundary action in Eq.˜19.
In this limit, referring to Fig.˜2, each boundary plaquette (green) has one edge in the bulk (black) with $A_{ij}=\theta_{j}-\theta_{i}$, and two edges straddling between the bulk and vacuum (green) which remain dynamical degrees of freedom. Substituting this into Eq.˜20, we can recombine terms into the gauge-invariant variables
@@ 2.3.1 Boundary XY Model at Infinite $\kappa$
$S^{\text{bdry}}_{\kappa\to\infty}=-\beta\sum_{\mathclap{\langle ij\rangle\in\partial X}}\cos(\vartheta_{j}-\vartheta_{i}),$ (22)
which is a 3D XY model at inverse temperature $\beta$. This must exhibit a continuous phase transition from a paramagnet at small $\beta$ to a spontaneously broken phase at large $\beta$. Thus we infer that along the $\kappa=\infty$ line in the phase diagram there is a boundary phase transition in the 3D XY universality class indicated at the top of Fig.˜1. We note that in the case of the $\mathbbm{Z}_{2}$ Abelian-Higgs model, the same mechanism generates a boundary Ising model at $\kappa=\infty$ [22].
2.3.2 Boundary Phase Transition at Finite $\kappa$
@@
These arguments are suggestive of the picture illustrated in Fig.˜1, with a boundary phase transition between the Higgs and confinement regimes of the bulk phase diagram. To test this assertion, we have carried out Monte Carlo simulations of the full 4D lattice gauge theory with boundary. We compute the local XY order parameter $\langle\vartheta_{i}\rangle$ on the boundary as well as gauge invariant bulk observables $\langle\Lambda_{\ell}\rangle$ and $\langle W_{p}\rangle$. The results are summarized in Fig.˜3. We find clear signs of a boundary phase transition, with Fig.˜3(a) showing the boundary order parameter behavior as a function of $\beta$ while holding $\kappa$ fixed showing behavior consistent with a continuous boundary phase transition. Fig.˜3(b) shows the Binder cumulant for the order parameter taken along a cut at $\kappa=2$ for different system sizes ranging from $L=16$ to $L=32$, showing crossing behavior consistent with a second-order transition. Lastly, Fig.˜3(c) shows the Binder parameter with $\beta$ scaled by $L^{1/\nu}$ using the 3D XY critical exponent $\nu\approx 0.67$ [36], showing excellent scaling collapse, confirming a second-order phase transition on the boundary even for only moderately large $\kappa$.
Monte Carlo simulations of the 3D XY model in the literature put the critical point at $\beta_{c}=0.45420(2)$ [36]. We find that $\beta_{c}$ tends towards this value in the large $\kappa$ limit. On general grounds we should expect that bulk fluctuations will serve to disorder the boundary. Therefore, by lowering $\kappa$ we expect the transition shifts to larger $\beta$ (lower effective temperature in the statistical model). This is indeed what we observe numerically. For even smaller $\kappa$ we find that the boundary transition line appears to intercept the bulk critical endpoint, as shown in Fig.˜1.
2.3.3 Tuning the Boundary Coupling
We now consider tuning the boundary coupling in Eq.˜19 relative to the bulk, parameterized by the dimensionless ratio
@@ 2.3.3 Tuning the Boundary Coupling
Indeed, this is precisely what we find numerically with the resulting transition lines shown in Fig.˜4. We find, for all $\alpha$, that the transition is present and that it drifts to larger $\beta$ as $\kappa$ is reduced, consistent with having to lower the temperature more to suppress the enhanced matter field fluctuations. For $\alpha<1$ the line shifts to larger values of $\beta$, and appears to separate from the tricritical point, instead terminating on the first order transition line separating the Higgs phase from the deconfined phase. For $\alpha>1$, the boundary transition line shifts to smaller values of $\beta$, and appears to continue to terminate on the tricritical point.
@@
2.4 Summary and Discussion: Abelian Case
In this section we have explored the case of the U(1) Abelian-Higgs model in $D$ spacetime dimensions with open boundary conditions of the form shown in Fig.˜2, governed by the action defined by Eqs.˜8, 19 and 18. In the case of open boundaries there is a global bulk charge symmetry for this model, because charge is not allowed to enter or leave the system. By the Gauss law, this is physically equivalent to a symmetry acting on the boundary of the system, yielding conservation of total electric flux through the boundary. We investigated the possibility of spontaneously breaking this symmetry.
In the $\cramped{\kappa\to\infty}$ limit, the bulk degrees of freedom are fully frozen while the boundary degrees of freedom remain fluctuating. The boundary theory in this limit can be written in the gauge-invariant form of a (D1)(D-1)-dimensional 0-form U(1) model, i.e. an XY model, Eq.˜22, which therefore exhibits a boundary phase transition at a critical value of $\beta$ in the appropriate XY universality class. By performing explicit numerical simulation in $D=4$ Euclidean spacetime dimensions and measuring the gauge-invariant order parameter, Fig.˜3, we found that this boundary phase transition persists away from the $\cramped{\kappa\to\infty}$ limit, and appears to stay in the 3D XY universality class. The transition line appears to terminate at a critical point in the bulk phase diagram and can be tuned by adjusting the boundary coupling relative to the bulk coupling, as shown in Fig.˜4.
This boundary transition line appears to delineate between the Higgs and confining regimes of the phase diagram. The idea that the Higgs phase can be characterized by boundary symmetry breaking was first raised in [22, 23]. This gives physical meaning to the non-gauge-invariant adage that the Higgs regime is a charge condensate by imposing open boundaries that make the symmetry physical. We note, however, that by tuning the boundary coupling relative to the bulk coupling, the location of the transition line in the phase diagram can be moved. This raises questions as to what the boundary phase transition implies as a probe of bulk physics, which we defer to the discussion at the end of the paper (Section 5.2). First we broaden our perspective by studying more complex non-Abelian models exhibiting similar physics in Section˜3.
A few comments on extensions and exceptional cases are in order before we proceed. These results naturally carry over to the discrete Abelian gauge groups $\mathbbm{Z}_{N}$ by restricting the Higgs field $\theta_{i}$ and gauge field $A_{ij}$ to take discrete values in multiples of $2\pi/N$. The boundary theory Eq.˜22 then becomes an $N$-state clock model. The $\mathbbm{Z}_{2}$ case was discussed in [22] where the boundary theory was shown to be an Ising model. They also can be generalized to higher-form extensions of the Abelian-Higgs model, which we discuss further in Section˜4.
@@ 2.4 Summary and Discussion: Abelian Case
3 Boundary Symmetry Breaking in Non-Abelian Higgs Models
@@
3.1 Preliminaries: Lattice Formulation
@@
where $g_{i}\in\mathcal{G}$ are arbitrary group elements associated to each site $i$. The minimal gauge-invariant quantity is the Wilson plaquette-loop (compare to Eq.˜6),
$W_{p}=\frac{1}{\mathrm{dim}({r})}\,\mathrm{Tr}_{\,{r}}\Big{[}\,\,{\prod_{\mathclap{\ell\in\partial p}}}^{P}U_{\ell}\Big{]}$ (27)
where the superscript $P$ on the product indicates path-ordering, and the trace may be taken in a representation r{r}. Normalizing by the dimension of the representation ensures that the trivial Wilson loop has unit magnitude. We focus primarily on the cases $\mathrm{SU}(N)$ and $\mathrm{SO}(N)$, taking the trace in the fundamental representation as $N\times N$ matrices.
3.1.1 Higgs Fields
For the Higgs field, many different models can be considered by putting the Higgs field in different representations of the gauge group. We consider two different types here for concreteness: vector-valued (fundamental representation) Higgs, and group-valued Higgs. These are both possible extensions of the Abelian U(1) rotor model considered in Section˜2, since a rotor may be viewed either as a fixed-length vector, or as a U(1) group element. In either case, the action is given by the Fradkin-Shenker form, Eq.˜7, the only difference being the definition of the Wilson link $\Lambda_{\ell}$. The generating function for the quantum theory is given by the Euclidean path integral, where the integration over the group-valued variables is performed with respect to the Haar measure.
The familiar model is the fundamental-Higgs, where the Higgs field is an $N$-component vector, as in the Standard Model and analogous to Eq.˜1. Denoting the Higgs vector at site $i$ by $\phi_{i}\in V_{i}$, we freeze the radial mode as in the Abelian Fradkin-Shenker model. Gauge transformations rotate the Higgs field as $\phi_{i}\to g_{i}^{\text{f}}\phi_{i}$ where $g_{i}^{\text{f}}$ is a group element in the fundamental matrix representation. The group-valued link variables define parallel-transport maps for the Higgs field, $U_{ij}:V_{j}\to V_{i}$, i.e. they related the color frames at neighboring sites, and the generalization of the gauge-invariant Wilson link observable, Eq.˜5, is